# RoPINN: Region Optimized Physics-Informed Neural Networks

### v1

@@ 5 Conclusion
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]&\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}\end{split}$ (35)
Suppose that at the first $K$ steps $\mathbb{E}[\delta_{t_{0}+K}]\leq\frac{2L}{\beta}$, Accumulating the above in equations recursively, we have the error accumulated to the first $K$ steps as follows:
@@
Then, following a similar process as Theorem 3.3(2), we can obtain:
@@
According to Lemma A.3, we have the gradient update operator $G_{\alpha,\boldsymbol{x}}^{\mathrm{region},x_{j}}$ for $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}$ satisfies the following inequality:
$G_{\alpha,\boldsymbol{x}}^{\mathrm{region},x_{j}}(\theta_{1})-G_{\alpha,\boldsymbol{x}}^{\mathrm{region},x_{j}}(\theta_{2})\leq(1+2\alpha\beta)\|\theta_{1}-\theta_{2}\|.$ (41)
Let $M=\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|}$, since $\alpha\leq\frac{1}{2\beta t}$, we can rewrite the Eq. (34) as follows:
$\begin{split}&\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\left((1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta}(1-M)\right).\\ \end{split}$ (42)
@@
As presented in Eq. (8) and (44), we approximate the region optimization with the Monte Carlo sampling method. For better efficiency, we propose to only sample one point at each iteration. However, this will cause an estimation error formalized in Theorem 3.9, which can be directly derived by the definition of standard deviation.
@@
$\begin{split}&\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\|\\ &=\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}}+\alpha_{t-1}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2}),\boldsymbol{z}_{2})\|\\ &=\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})+\alpha_{t-1}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\nabla^{2}_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{x}^{\prime})\|\\ &\leq\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\|+\alpha_{t-1}L\beta.\end{split}$ (48)
@@
$\begin{split}\lim_{t\to\infty}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},z_{i})&=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{i})\\ \lim_{t\to\infty}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},z_{i})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-j+1}},z_{j})&=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{i})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{j}).\end{split}$ (49)

### v2

@@ 5 Conclusion
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]&\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|},\end{split}$ (39)
where we do not consider the benefits brought by the overlap area of $\boldsymbol{x}+\Omega_{r}$ and $\boldsymbol{x}^{\prime}+\Omega_{r}$.
Suppose that at the first $K$ steps $\mathbb{E}[\delta_{t_{0}+K}]\leq\frac{2L}{\beta}$, Accumulating the above in equations recursively, we have the generalization error bound accumulated to the first $K$ steps as follows:
@@
Then, considering the all $T$ steps, we have
$\begin{split}\mathbb{E}\left[\|\theta_{T}-\theta_{T}^{\prime}\||\delta_{t_{0}}=0\right]&\leq\Delta+\sum_{t=t_{0}+K+1}^{T}\left\{\Pi_{k=t+1}^{T}\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\right\}\frac{2L}{\beta t|\mathcal{S}|}\\ &\leq\sum_{t=t_{0}+1}^{T}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2L}{\beta t|\mathcal{S}|}-JM^{2}\quad\quad\quad\text{($\sum_{k=t+1}^{T}\frac{1}{k}\leq\log\frac{T}{t}$)}\\ &=\frac{2L}{\beta|\mathcal{S}|}T^{1-\frac{1}{|\mathcal{S}|}}\sum_{t=t_{0}+1}^{T}t^{-(1-\frac{1}{|\mathcal{S}|})-1}-JM^{2}\\ &\leq\frac{2L}{\beta|\mathcal{S}|}T^{1-\frac{1}{|\mathcal{S}|}}\frac{1}{1-\frac{1}{|\mathcal{S}|}}\left(t_{0}^{-(1-\frac{1}{|\mathcal{S}|})}-T^{-(1-\frac{1}{|\mathcal{S}|})}\right)-JM^{2}.\quad\text{(Integral approximation)}\\ \end{split}$ (41)
Thus, following a similar proof process as Theorem 3.3(2), we can obtain:
@@
According to Lemma A.3, we have the gradient update operator $G_{\alpha_{t},\boldsymbol{x}}^{\mathrm{region},x_{j}}$ for $\frac{\partial}{\partial x_{j}}\mathcal{L}_{r}^{\mathrm{region}}$ satisfies the following inequality:
$\|G_{\alpha_{t},\boldsymbol{x}}^{\mathrm{region},x_{j}}(\theta_{1})-G_{\alpha_{t},\boldsymbol{x}}^{\mathrm{region},x_{j}}(\theta_{2})\|\leq(1+2\alpha_{t}\beta)\|\theta_{1}-\theta_{2}\|.$ (46)
Let $M=\frac{|\Omega_{r}|}{|\Omega|}$, since $\alpha_{t}\leq\frac{1}{2\beta t}$, we can rewrite the Eq. (37) as follows:
$\begin{split}&\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\left((1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta}(1-M)\right),\\ \end{split}$ (47)
where the second term is derived from Eq. (38) by substituting $L$ to $2L$ and $\beta$ to $2\beta$, which is:
$\begin{split}\mathbb{E}\left[\|G^{\mathrm{region},x_{j}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region},x_{j}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]&\leq(1+\frac{2\alpha_{t}\beta|\Omega_{r}|}{|\Omega|})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+4\alpha_{t}L(1-\frac{|\Omega_{r}|}{|\Omega|})\\ &=(1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta t}(1-M).\end{split}$ (48)
@@
As presented in Eq. (8) and (50), we approximate the region optimization with the Monte Carlo sampling method. For better efficiency, we propose only to sample one point at each iteration. However, this will cause an estimation error formalized in Theorem 3.9, which can be directly derived by the definition of standard deviation as follows:
Proof.
According to the definition of $\mathcal{L}_{r}^{\mathrm{region}}$ in Eq. (5), we get
$\begin{split}&\mathbb{E}_{\boldsymbol{\xi}\sim U(\Omega_{r})}\left[\left\|\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})-\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})\right\|^{2}\right]^{\frac{1}{2}}\\ &=\mathbb{E}_{\boldsymbol{\xi}\sim U(\Omega_{r})}\left[\left\|\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})-\nabla_{\theta}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right\|^{2}\right]^{\frac{1}{2}}\\ &=\left\|\sigma_{\boldsymbol{\xi}\sim U(\Omega_{r})}\left(\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\right)\right\|.\end{split}$ (54)
@@
$\begin{split}&\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\\ &=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}+\alpha_{t-1}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})},\boldsymbol{z}_{2})\\ &=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})+\alpha_{t-1}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\nabla^{2}_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{x}^{\prime}).\end{split}$ (55)
Thus, the following inequality holds:
$\begin{split}&{\big{|}}\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\right\|-\left\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\right\|{\big{|}}\\ &\leq\left\|\left(\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\right)-\left(\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{1})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\boldsymbol{z}_{2})\right)\right\|\\ &=\left\|\alpha_{t-1}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{z}_{2})\nabla^{2}_{\theta}\mathcal{L}(u_{\theta_{t-1}},\boldsymbol{x}^{\prime})\right\|\\ &\leq\beta L\alpha_{t-1}.\end{split}$ (56)
@@
$\begin{split}\lim_{t\to\infty}\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},z_{i})&=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{i})\\ \lim_{t\to\infty}\left(\nabla_{\theta}\mathcal{L}(u_{\theta_{t-i+1}},z_{i})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t-j+1}},z_{j})\right)&=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{i})-\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},z_{j}).\end{split}$ (57)
@@
B.4 Proof for Region Optimization with Gradient Estimation Error (Theorem 3.12)
Convex setting
Firstly, we would like to prove the convex case as follows.
Proof.
Similar to the proof of region optimization in Appendix A.3, at the $t$-th step, we can obtain the following equation:
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\right]&=(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t}^{\prime})\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right],\end{split}$ (60)
where $G^{\mathrm{approx}}_{\alpha,\boldsymbol{x}}(\theta)=\theta-\alpha\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{approx}}(\theta,\boldsymbol{x})=\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta,\boldsymbol{x}+\boldsymbol{\xi}),\boldsymbol{\xi}\sim U(\Omega_{r})$. Suppose that we have sampled ${\boldsymbol{\xi},\boldsymbol{\xi}^{\prime}\in\Omega_{r}}$, the second term on the right part can be bounded as follows:
[…]
$\begin{split}\mathcal{E}_{\mathrm{gen}}\leq\left(L(1-\frac{|\Omega_{r}|}{|\Omega|})+{{\mathcal{E}_{r,\mathrm{grad}}}}\right)\frac{2L}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}.\end{split}$ (62)
Non-convex setting
Similarly, we can prove the non-convex setting as follows.
Proof.
It is easy to prove that $\mathcal{L}_{r}^{\mathrm{approx}}(\theta,\boldsymbol{x})$ is still $L$-Lipchitz-$\beta$-smoothness for $\theta$. For clarity, we define that $M=\frac{|\Omega_{r}|}{|\Omega|}$. Thus, based on Eq. (38), we have the following derivations.
If $\mathbb{E}(\delta_{t})\leq\frac{2L}{\beta}-\frac{2}{\beta M}\mathcal{E}_{r,\mathrm{grad}}$, we have:
$\begin{split}&\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{approx}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]\\ &+\frac{1}{|\mathcal{S}|}\left(\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\right)\\ &+\frac{1}{|\mathcal{S}|}\left(\mathbb{E}\left[\|\alpha_{t}\nabla_{\theta}\mathcal{L}(\theta_{t},\boldsymbol{x}+\boldsymbol{\xi})-\alpha_{t}\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(\theta_{t},\boldsymbol{x})\|\right]\right)\\ &+\frac{1}{|\mathcal{S}|}\left(\mathbb{E}\left[\|\alpha_{t}\nabla_{\theta}\mathcal{L}(\theta_{t},\boldsymbol{x}^{\prime}+\boldsymbol{\xi}^{\prime})-\alpha_{t}\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(\theta_{t}^{\prime},\boldsymbol{x}^{\prime})\|\right]\right)\\ &\leq(1+\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|})\mathbb{E}[\delta_{t}]+\frac{2\alpha_{t}}{|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right)\\ &\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2}{\beta t|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right).\end{split}$ (63)
Otherwise, we still consider the following inequality:
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]&\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|},\end{split}$ (64)
Suppose that at the first $K^{\prime}$ steps $\mathbb{E}[\delta_{t_{0}+K^{\prime}}]\leq\frac{2L}{\beta}-\frac{2}{\beta M}\mathcal{E}_{r,\mathrm{grad}}$.
Accumulating the above in equations recursively, we have the generalization error bound accumulated to the first $K^{\prime}$ steps as follows:
$\begin{split}\Delta&=\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\left\{\Pi_{k=t+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left(\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|}\right)\Pi_{k=t_{0}+K^{\prime}+1}^{T}\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\right\}\frac{2}{\beta t|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right)\\ &\leq\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t_{0}+K^{\prime}}+(1-\frac{1-M}{|\mathcal{S}|})\log\frac{t_{0}+K^{\prime}}{t}\right)\frac{2}{\beta t|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right)\\ &=\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}+\frac{M}{|\mathcal{S}|}\log\frac{t_{0}+K^{\prime}}{t}\right)\frac{2}{\beta t|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right)\\ &\leq\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2}{\beta t|\mathcal{S}|}\left(L(1-M)+{\mathcal{E}_{r,\mathrm{grad}}}\right)(\frac{t_{0}+K^{\prime}}{t})^{\frac{M}{|\mathcal{S}|}}\\ &\leq\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2L}{\beta t|\mathcal{S}|}-\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2}{\beta t|\mathcal{S}|}\left(LM^{2}+\mathcal{E}_{r,\mathrm{grad}}(1+M)\right)\\ &=\sum_{t=t_{0}+1}^{t_{0}+K^{\prime}}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2L}{\beta t|\mathcal{S}|}-J^{\prime}LM^{2}+J^{\prime}\mathcal{E}_{r,\mathrm{grad}}(1+M),\end{split}$ (65)
where $J^{\prime}$ is a finite value that depends on the training property of beginning iterations, namely $K^{\prime}$ and $t_{0}$. The last inequality is from $(\frac{t_{0}+K^{\prime}}{t})^{\frac{M}{|\mathcal{S}|}}\leq(1+M)$, when $|\mathcal{S}|$ is sufficient enough.
Then, considering the all $T$ steps, we have
$\begin{split}&\mathbb{E}\left[\|\theta_{T}-\theta_{T}^{\prime}\||\delta_{t_{0}}=0\right]\\ &\leq\Delta+\sum_{t=t_{0}+K+1}^{T}\left\{\Pi_{k=t+1}^{T}\mathrm{exp}\left(\frac{1}{t}-\frac{1}{t|\mathcal{S}|}\right)\right\}\frac{2L}{\beta t|\mathcal{S}|}\\ &\leq\sum_{t=t_{0}+1}^{T}\mathrm{exp}\left((1-\frac{1}{|\mathcal{S}|})\log\frac{T}{t}\right)\frac{2L}{\beta t|\mathcal{S}|}-J^{\prime}M^{2}+J^{\prime}\mathcal{E}_{r,\mathrm{grad}}(1+M)\quad\quad\quad\text{($\sum_{k=t+1}^{T}\frac{1}{k}\leq\log\frac{T}{t}$)}\\ &=\frac{2L}{\beta|\mathcal{S}|}T^{1-\frac{1}{|\mathcal{S}|}}\sum_{t=t_{0}+1}^{T}t^{-(1-\frac{1}{|\mathcal{S}|})-1}-J^{\prime}M^{2}+J^{\prime}\mathcal{E}_{r,\mathrm{grad}}(1+M)\\ &\leq\frac{2L}{\beta|\mathcal{S}|}T^{1-\frac{1}{|\mathcal{S}|}}\frac{1}{1-\frac{1}{|\mathcal{S}|}}\left(t_{0}^{-(1-\frac{1}{|\mathcal{S}|})}-T^{-(1-\frac{1}{|\mathcal{S}|})}\right)-J^{\prime}M^{2}+J^{\prime}\mathcal{E}_{r,\mathrm{grad}}(1+M).\quad\text{(Integral approximation)}\\ \end{split}$ (66)
Next, following the proof in Appendix A.3, we can obtain the generalization bound as follows:
$\mathbb{E}\left[|\mathcal{L}({u_{\theta_{T}}},\boldsymbol{x})-\mathcal{L}({u_{\theta_{T}^{\prime}}},\boldsymbol{x})|\right]\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-J^{\prime}LM^{2}+J^{\prime}\mathcal{E}_{r,\mathrm{grad}}(1+M).$ (67)
Note that $K^{\prime}$ does not exist when $\frac{2L}{\beta}<\frac{2}{\beta M}\mathcal{E}_{r,\mathrm{grad}}$, then $J^{\prime}=0$, which corresponds to the situation that the region size is too large and brings serious gradient estimation error. Introducing “region” cannot bring a better generalization bound in this case. ∎